The velocity distribution in a two dimensional steady flow field is xy plan is v=(Ax-B)i+(C-Ay)j,A=2s-1,B=5ms-1,C=3ms-1. The coordinates are measured in meters and body force distributing is gx=-gk.Does the velocity field represent an incompressible fluid?Find the stagnation point of the flow field. Obtain expression for the pressures gradient in the flow field. Evaluate the difference at(x,y)=(1,3) and origin if density is 1.2kg/m3

1
Expert's answer
2021-07-13T05:16:33-0400

Solution;

a) Incompressible flow

"\frac{du}{dx}+\frac{dv}{dy}=0"

Given:

u=Ax-B

"\frac{du}{dx}"=A

v=C-Ay

"\frac{dv}{dy}" =-A

"\frac{du}{dx}+\frac{dv}{dx}" =A+-A=0

Answer

The velocity field represents an incompressible flow.

b) Stagnation point of the flow field;

At stagnation point ;

"\overrightarrow{V}" =0

Since

"\overrightarrow{V}" =(Ax-B)i+(C-Ay)j

And A=2,B=5 and C=3

u=2x-5=0

x=5/2

v=3-2y

y=3/2

Answer

Stagnation point

"(\frac52,\frac32)"

c) Expression for the pressure gradient.

From Euler's equation of ideal flow,

Assuming a steady flow;

"\rho" gx-∆P="\rho" ["u\frac{dv}{dx}+v\frac{dv}{dy}" ]

From the given data

"-\rho g_xk-∆P=\rho[(Ax-B)Ai+(C-Ay)-Aj]"

Replace the values of A B and C

"-\rho g_xk-∆P=\rho[(4x-10)i+(4y-6)j]"

Which can be rewritten as;

"∆P=-\rho[(4x-10)i+(4y-6)j+g_xk]"

d) Evaluate the difference at (1,3) and Origin(0,0) if "\rho=1.2kg/m^3"

P1,3-P0,0="-\rho(\int_0^1(4x-10)dx+\int_0^3(4y-6)dy+\int_0^0gdz)"

P(1,3)-P(0,0)="-\rho[(2x^2-10x)|_0^1+(2y^2-6y)|_0^3]"

P(1,3)-P(0,0)=-1.2(-8)=9.6N/m2

Answer

9.6N/m2